OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (jackellyn):

If two line are parallel they have the same slope The form y=mx+b gives the slope (m) and the y intercept (b). The point (5,3) has a y intercept of +3. y=mx +3 What is the slope, or m?

OpenStudy (anonymous):

I got y = 2x − 7

OpenStudy (jackellyn):

Yep you're right, after putting it in point slope form you get y-3 = 2(x1 -5) y-3 = 2x -10 y = 2x -7